OpenStudy (anonymous):

8 over 9 equals 4 over 5 with k next to 4 over 5 what is k?

OpenStudy (anonymous):

8/9 = 4k/5, cross multiply and set them equal to each other...40 = 32k, k = 40/32 or 5/4

OpenStudy (anonymous):

oop sry it's 40 = 36K

OpenStudy (anonymous):

so K = 10/9

OpenStudy (anonymous):

the question is 8/9 equals 4/5 k what is k?

OpenStudy (anonymous):

yea i know

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

np